complete the missing? (_5^11)B + ______ → (_6^13)C + (_1^1)H
1
Expert's answer
2019-11-22T05:34:49-0500

Let us denote the unknown element by X. We assume it's atomic number to be Z and it's Atomic mass to be A.


"_5B^{11} +_ZX^{A} \to _6C^{13}+_1H^1"


Now, equating values for Z and A we get;

"11+A=13+1"

"\implies A=14-11"

"=3"

Similarly; "5+Z=6+1"

"\implies Z=2"


As, Atomic number is 2, clearly the element is Helium.

And as it's Atomic mass is 3, it has 1 neutron.

Thus the element "_ZX^{A}" can be identified as "_2He^{3}" , which the Helium-3 isotope of Helium.